OpenStudy (anonymous):

find the sum and product of the roots X^2-7x+2=0

OpenStudy (zzr0ck3r):

what are the roots?

OpenStudy (zzr0ck3r):

\[x^2-7x+2=0\\(x-\frac{7}{2})^2=-2+(\frac{7}{2})^2\\(x-\frac{7}{2})^2=\frac{-8}{4}+\frac{49}{4}\\(x-\frac{7}{2})^2=\frac{41}{4}\\x-\frac{7}{2}=\pm \sqrt{\frac{41}{4}}\\x=\pm \sqrt{\frac{41}{4}}+\frac{7}{2}\\x=\pm\frac{\sqrt{41}+7}{2}\\so\\x=\frac{\sqrt{41}+7}{2}\\and\\x=\frac{7-\sqrt{41}}{2}\]

OpenStudy (zzr0ck3r):

\[\frac{-(\sqrt{41}+7)}{2}\]

OpenStudy (zzr0ck3r):

that's what the last line should say, but we are not done. do you know the answer?

OpenStudy (zzr0ck3r):

we have not answered either question yet. Tell me what you think the question is trying to get you to do? because im not going to give you the final answer, you need to figure it out. I have don't most of the work for you.

OpenStudy (zzr0ck3r):

1st) we need to find the roots 2nd) we need to add the roots to each other 3rd) we need to multiply the roots with each other my question to you, given what I have shown you what are the roots?

OpenStudy (anonymous):

going with 1,0 as the roots thx for your time

OpenStudy (anonymous):

close off or just need to rethink the problem

OpenStudy (zzr0ck3r):

I already told you the roots....np gl